Two forces of 12N and 9N are acting at a point . The ratio of the magnitude of maximum and minimum resultant is.......

Grade:11

4 Answers

Arun
6 years ago
The resultant will be maximum when they act in same direction hence making an angle of 0 degree.
Then maximum = 12+9= 21 N
 
Now, resultant will be minimum if they act in opposite direction i.e. making an angle of 180 degree.
Then minimum= 12-9= 3 N
Thus ratio of magnitude of maximum and minimum
resultant = 21/3= 7
Shailendra Kumar Sharma
6 years ago
Resultant of two forces p and q is given by R2= P2+q2+ 2pqCos(theta) where there is the angle between the forces Now maximum value of cos(theta) will be 1 resulting in R2= P2+q2+ 2pqAnd R = p+q So maximum resultant will be 12+9=21About minimum value of Cos(theta) = -1Resulting in R = p-q =3So the ratio of maximum to minimum will be 7
PRABHAT
5 years ago
Yup , u can use pallelogram law of vector addition.To get resulting force(magnitude), use formula ,( a^2 + b^2 + 2(a)(b)(Cosx) )^1/2a & b are two vectors having “x” angle between them.So use this to get resulting force(magnitude)( 10^2 + 8^2 + 2(10)(8)cos60° )^1/2, Solve and get your answer ready!!!And you can also get the direction , again from pallelogram law .
PRABHAT
5 years ago
Answered by Prabhat Kumar Yadav 
Class 9
 
 The resultant of two forces acting on the same point is maximum when the forces are acting in the same direction and minimum when they operate in opposite direction.
When the two forces are acting in the same direction the angle between the direction to the two forces is 0. In this case the resultant force is sum of the two forces and its direction is same as the two independent forces.
In the question the magnitude of the two forces are given as 12 N and 9 N.
Therefor their net magnitude will be 12 + 9 = 21 N
Again 12-9 =3
Ratio 3:21 
 
So Ans :-- 1:7